OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (10, 35) and (150, 55), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

OpenStudy (anonymous):

1/7 Is what I got, but I may be wrong.

OpenStudy (anonymous):

hi

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( 55- 35)/(150-10) m = 20/140 m = 1/7 So the slope of the line that passes through the two points is 1/7

OpenStudy (anonymous):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

;)